Martha’s Vineyard Marine Supply is a wholesaler for a large variety of boating and fishing equipment. The company’s controller, Mathew Knight, has recently completed a cost study of the firm’s material-handling department in which he used work measurement to quantify the department’s activity. The control factor unit used in the work-measurement study was hundreds of pounds of equipment unloaded or loaded at the company’s loading dock. Knight compiled the following data.

Units of Activity
Month (hundreds of pounds of
equipment loaded or unloaded)
Material-Handling
Department Costs
January 3,000 $ 12,900
February 2,800 12,500
March 2,500 11,850
April 2,200 11,400
May 3,400 12,300
June 3,600 14,350
July 3,200 13,200
August 3,000 12,600
September 3,800 13,560
October 2,300 11,650
November 2,400 12,550
December 2,600 12,550
  1. Predict the firm’s material-handling department’s costs for a month when 2,000 units of activity are recorded. (Do not round intermediate calculations. Round variable cost per unit to 2 decimal places and final answer to the nearest whole dollar amount.)
Answer #1

Using High-Low Method

High-Low Method

Units of Activity

Material-Handling
Department Costs

High activity level

3800

13560

Low activity level

2200

11400

Change

1600

2160

Variable cost = (cost of highest activity – cost of lowest activity)/(highest activity – lowest activity) = (13560-11400)/(3800-2000) = 2160/1600 =$1.35 per unit

Fixed costs = total cost – total variable cost = 13560-(3800*1.35) = $8430

Cost function: Y=a+bX

Therefore,

Y = $8430 + $1.35X

Now,

firm’s material-handling department’s costs for a month when 2,000 units of activity are recorded = 8430+(2000*1.35) = $11130